# A Theoretical Perspective on Why Stochastic Population Update Needs an Archive in Evolutionary Multi-objective Optimization

### v2

@@ 1 Introduction
SMS-EMOA $O(\mu n^{k}\cdot\min\{1,\sqrt{k}\mu/2^{k/2}\})$ Bian et al. (2024b)
@@
NSGA-II $O(\mu\sqrt{k}(n/2)^{k})$ Bian et al. (2024b)
@@
Very recently, Bian et al. Bian et al. (2024b) analytically showed that introducing randomness in the population update process of MOEAs (called stochastic population update, SPU) can help the search. Specifically, the study proved that for SMS-EMOA solving the common benchmark problem OneJumpZeroJump, when $k=n/2-\Omega(n)$, using SPU can bring an acceleration of $\Omega(2^{k/2}/(\sqrt{k}\mu^{2}))=\Omega(2^{k/2}/(\sqrt{k}(n-2k+4)^{2}))$ on the expected running time, where $n$ denotes the problem size, $k$ ($2\leq k<n/2$) denotes the parameter of OneJumpZeroJump, and $\mu$ denotes the population size. Subsequently, Zheng and Doerr Zheng and Doerr (2024b) extended SMS-EMOA to solve a many-objective problem, mmOneJumpZeroJump, showing that the same SPU can bring an acceleration of $\Theta(2^{k}/\mu)=\Theta(2^{k}/(2n/m-2k+3)^{m/2})$ as well, where $m$ is the number of objectives. These works echoed the empirical studies that show the benefit of considering non-elitism in MOEAs Tanabe and Ishibuchi (2019); Liang et al. (2023a).
SPU in Bian et al. (2024b); Zheng and Doerr (2024b) introduces randomness by randomly selecting a proportion $p_{s}$ of the combined set of the current population and offspring solution(s) to be directly preserved into the next generation. This essentially gives a chance for inferior solutions to survive, which enables the evolutionary search to go along inferior regions which may be close to Pareto optimal regions. However, a cost of this method is that there is less space for the best solutions in the population. Some very best solutions to the problem (e.g., globally non-dominated solutions), even found by an MOEA, may be discarded during the population update process. This necessitates a large population used. Unfortunately, when the population size is large, the benefit of using SPU may vanish. This is because the benefit of SPU comes from the operation on inferior solutions while the large population size will lead to a small probability of selecting these solutions. For example, for SMS-EMOA solving OneJumpZeroJump, the acceleration of $\Omega(2^{k/2}/(\sqrt{k}\mu^{2}))$ brought by SPU Bian et al. (2024b) will vanish when the population size $\mu$ is exponential w.r.t. $k$, e.g., $k=\log n$; for SMS-EMOA solving mmOneJumpZeroJump, the acceleration of $\Theta(2^{k}/(2n/m-2k+3)^{m/2})$ Zheng and Doerr (2024b) will vanish when the number $m$ of objectives is large, e.g., $m\geq k$, because the population size $\mu=(2n/m-2k+3)^{m/2}$ increases rapidly with $m$. This dilemma impacts both the effectiveness and practicality of using SPU in MOEAs.
@@
  • We theoretically show that incorporating an archive mechanism with SPU can reduce the upper bound on the expected running time (even exponentially). For example, comparing Theorems 1 and 3 in Table 1, the expected running time of SMS-EMOA with SPU for solving OneJumpZeroJump, no matter whether an archive is used or not, is $O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$, where $C=e\mu/(p_{s}(1-p_{c}))$. The key difference is that using an archive allows a constant population size, resulting in a significantly smaller $C$ and thus reducing the upper bound significantly. Note that Bian et al. Bian et al. (2024a) recently proved the effectiveness of using an archive (bringing polynomial acceleration) for MOEAs, while our analysis for MOEAs with SPU reveals that even exponential acceleration can be obtained.
@@
  • In addition, our derived running time bounds for MOEAs with SPU in the second column of Table 1 are significantly better than the previously known ones Bian et al. (2024b); Zheng and Doerr (2024b). This improvement stems from the analysis method of considering both the number and size of jumps across gap of dominated solutions. The method happens to share similarities with that of proving Lemma 12 in doerr2024hyper. Moreover, our running time bounds are more general, as we consider variable survival probability $p_{s}$ and crossover probability $p_{c}$.
@@
Finally, we give a brief overview about the running time analysis of MOEAs. Over the last decade, there has been an increasing interest for the evolutionary theory community to study MOEAs. Early theoretical works Laumanns et al. (2004a, b); Neumann (2007); Giel and Lehre (2010); Neumann and Theile (2010); Doerr et al. (2013); Qian et al. (2013, 2016); Bian et al. (2018) mainly focus on analyzing the expected running time of a simple MOEA like GSEMO/SEMO. Recently, researchers have begun to examine practical MOEAs. Huang et al. Huang et al. (2021) investigated MOEA/D, assessing the effectiveness of different decomposition methods. Zheng et al. Zheng et al. (2022) conducted the first theoretical analysis of NSGA-II. Bian et al. Bian et al. (2024b) analyzed the running time of SMS-EMOA and showed that SPU can bring acceleration. Moreover, Wietheger and Doerr Wietheger and Doerr (2023) demonstrated that NSGA-III Deb and Jain (2014) exhibits superior performance over NSGA-II in solving the tri-objective problem 33OneMinMax. Ren et al. Ren et al. (2024a) analyzed the running time of SPEA2 on three $m$-objective problems. Some other works on well-established MOEAs include Bian and Qian (2022); Zheng and Doerr (2022, 2024a); Cerf et al. (2023); Dang et al. (2023a, b); Doerr and Qu (2023b, c); Doerr et al. (2024a, b); Lu et al. (2024); Opris et al. (2024); Opris (2024); Ren et al. (2024b).
@@ 3 Running Time Analysis of SMS-EMOA
In this section, we prove that using an archive can bring (even exponential) speedup for the well-established MOEA, SMS-EMOA Beume et al. (2007), with stochastic population update (SPU) solving the OneJumpZeroJump problem. In Section 3.1, we first introduce SMS-EMOA with SPU, and its expected running time for solving OneJumpZeroJump, which is much tighter than the previous results Bian et al. (2024b); Zheng and Doerr (2024b). Then, in Section 3.2, we prove that using an archive can significantly reduce the upper bound on the expected running time. Due to space limitation, all mathematical proofs could only be sketched or had to be omitted completely.
@@ 3.1 SMS-EMOA with SPU
In Bian et al. (2024b), the SPU method is introduced and shown to be beneficial for the search of MOEAs. During population updates, SPU randomly selects a proportion $p_{s}$ of the current population and the offspring solution(s) to directly survive into the next generation and the removed part is selected from the rest solutions. This implies that each solution, including the worst solution in the population, has at least a probability $p_{s}$ of surviving to the next generation. Specifically, SPU of SMS-EMOA as presented in Algorithm 3 is used to replace the original Population Update procedure in line 12 of Algorithm 1. Note that $p_{s}$ is set to $1/2$ in Bian et al. (2024b), while we consider a general $p_{s}$ here.
@@
The expected running time of SMS-EMOA with SPU for solving OneJumpZeroJump has been proven to be $O(\mu n^{k}\cdot\min\{1,\sqrt{k}\mu/2^{k/2}\})$ Bian et al. (2024b), which is better than that, i.e., $O(\mu n^{k})$, of the original SMS-EMOA. Intuitively, by by introducing randomness into the population update, the evolutionary search has a chance to go along inferior regions which are close to Pareto optimal regions, thereby making the search easier. Here, we re-prove a tighter upper bound on the expected running time of SMS-EMOA with SPU for solving OneJumpZeroJump, as shown in Theorem 1. It is also more general, as it considers a survival probability $p_{s}\in[1/(\mu+1),1-o(1/\mu))$, rather than just $p_{s}=1/2$ as in Bian et al. (2024b). Note that the running time of EAs is often measured by the number of fitness evaluations.
@@
Since the objective vectors in $F_{I}^{*}$ will always be maintained by Lemma 2, we could use the conclusion of Theorem 1 in Bian et al. (2024b), which proves that the expected number of fitness evaluations of the first phase is $O(\mu(n\log n+k^{k}))$, where the term $O(\mu k^{k})$ is the expected number of fitness evaluations for finding one objective vector in $F_{I}^{*}$ when event $E$, i.e., any solution in the initial population has at most (k1)(k-1) or at least (nk+1)(n-k+1) 1-bits, happens. By Chernoff bound and $n-2k=\Omega(n)$, an initial solution has at most (k1)(k-1) or at least (nk+1)(n-k+1) 1-bits with probability $\exp(-\Omega(n))$, implying that event $E$ happens with probability $\exp(-\Omega(n))^{\mu}$. Thus, the term $O(\mu k^{k})$ can actually be omitted, implying that the expected number of fitness evaluations of the first phase is $O(\mu n\log n)$.
@@
Under the same conditions as the previous results $O(\mu n^{k}\cdot\min\{1,\sqrt{k}\mu/2^{k/2}\})$ in Bian et al. (2024b) and $O(\mu n^{k}\cdot\min\{1,\mu/2^{k}\})$ in Zheng and Doerr (2024b), where SPU is used with $p_{s}=1/2$ and crossover probability $p_{c}=0$, our bound in Theorem 1 becomes $O(\mu n^{k}\cdot(e\ln(2e\mu)/k)^{k-1})$. Since the bound in Zheng and Doerr (2024b) is tighter than that in Bian et al. (2024b), we focus on comparing our result only with Zheng and Doerr (2024b). When $k>2e\ln(2e\mu)$, our bound $O(\mu n^{k}\cdot(e\ln(2e\mu)/k)^{k-1})$ brings an improvement ratio of $\Theta(\mu(k/(2e\ln(2e\mu)))^{k-1})$, which can be exponential when $k$ is large, e.g., $k=n/8$. For a very small range $\log\mu\leq k<2e\ln(2e\mu)$, our bound shows no advantage, which is because to minimize Eq. (6), we choose $M=\lceil(k-1)/\ln C-1\rceil$ to round up to the nearest integer, leading to an over-relaxation. If we instead set $M=1$, the same bound can be obtained. The reason of our better bound is in the second phase of finding the Pareto optimal solution $1^{n}$: 1) We consider $M+1$ “jumps” across the gap between dominated solutions, and account for all possible jump sizes to find $1^{n}$; 2) We select the optimal number of jumps, i.e., $M=\lceil(k-1)/\ln C-1\rceil$. Note that Bian et al. Bian et al. (2024b) considered only two fixed-size jumps, and although Zheng and Doerr Zheng and Doerr (2024b) accounted for all possible jump sizes, they also reduced the process to two jumps.
@@ 3.2 An Archive is Provably Helpful
In the last section, we have proved that for SMS-EMOA with SPU solving OneJumpZeroJump, the expected running time is $O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$, where $C=e\mu/(p_{s}(1-p_{c}))$. From the analysis, we find that SPU benefits evolutionary search by exploring inferior regions that are close to Pareto optimal areas, but it also requires a larger population size $\mu\geq(n-2k+4)/(1-p_{s})$ to preserve the Pareto optimal solutions discovered. The greater the randomness introduced by SPU (i.e., the larger the value of $p_{s}$), the larger the population size needed. Note that a large population size may diminish the benefit of SPU, because it will lead to a very small probability of selecting specific dominated solutions for reproduction, which is required by SPU. For example, when the population size $\mu$ is exponential w.r.t. $k$, e.g., $k=\log n$, the improvement by SPU will vanish, compared to the expected running time $O(\mu n^{k})$ without SPU Bian et al. (2024b).
@@
The smaller population size allows a larger probability of selecting inferior solutions that are close to Pareto optimal areas for reproduction, thus leading to speedup. When $k$ is limited, using only SPU may not bring acceleration. For example, when $k=e\ln(8en)$, if using SPU with $p_{s}=1/2$, $\mu=2n$, and $p_{c}=1/2$, the expected running time of SMS-EMOA on OneJumpZeroJump is $O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})=O(n^{k+1})$ (where $C=e\mu/(p_{s}(1-p_{c}))=8en$), implying no acceleration compared to $O(\mu n^{k})=O(n^{k+1})$ without SPU Bian et al. (2024b). However, by adding an archive and reducing the population size $\mu$ to $5$, while keeping other settings unchanged, the bound reduces to $O(n^{k}\cdot(e\ln(20e)/k)^{k-1})$, implying a superpolynomial reduction in the upper bound as $k=e\ln(8en)$. Note that we set $\mu=5$ to satisfy the condition of Theorem 3, i.e., to make $p_{s}=1/2\in[1/(\mu+1),(\mu-2)/(\mu+1)]$. On the other hand, when $k$ is large, using SPU alone can lead to exponential acceleration, and the addition of an archive can further enhance this acceleration. For example, when $k=n/8$, if using SPU with $p_{s}=1/2$, $\mu=2n$, and $p_{c}=1/2$, the expected running time is $O(\mu n^{k}\cdot(8e\ln(8en)/n)^{n/8-1})$, which implies an exponential reduction in the upper bound compared to $O(\mu n^{k})$ without SPU. By adding an archive and setting $\mu$ to $5$, the bound reduces to $O(n^{k}\cdot(8e\ln(20e)/n)^{n/8-1})$, resulting in an improvement ratio of $\Theta(n(\ln(8en)/\ln(20e))^{n/8-1})$, i.e., an exponential reduction in the upper bound.
We also note that the recent work Bian et al. (2024a) has proven that using an archive can provide polynomial acceleration for MOEAs for the first time. For example, for SMS-EMOA solving OneMinMax, the expected running time is $O(\mu n\log n)$ both with and without an archive, but the archive allows for a constant $\mu$, achieving an acceleration of $\Theta(n)$. Our work gives another theoretical evidence for the effectiveness of using an archive, and further shows that superpolynomial or even exponential acceleration can be achieved. In our analysis with SPU, a sequence of $M+1$ consecutive jumps are required, which implies continuously selecting specific solutions for reproduction; using an archive allows a small population size, which increases the selection probability significantly and thus leads to greater acceleration.

### v3

@@ 1 Introduction
SMS-EMOA $O(\mu n^{k}\cdot\min\{1,\sqrt{k}\mu/2^{k/2}\})$ Bian et al. (2025)
@@
NSGA-II $O(\mu\sqrt{k}(n/2)^{k})$ Bian et al. (2025)
@@
Very recently, Bian et al. Bian et al. (2025) analytically showed that introducing randomness in the population update process of MOEAs (called stochastic population update, SPU) can help the search. Specifically, the study proved that for SMS-EMOA solving the common benchmark problem OneJumpZeroJump, when $k=n/2-\Omega(n)$, using SPU can bring an acceleration of $\Omega(2^{k/2}/(\sqrt{k}\mu^{2}))=\Omega(2^{k/2}/(\sqrt{k}(n-2k+4)^{2}))$ on the expected running time, where $n$ denotes the problem size, $k$ ($2\leq k<n/2$) denotes the parameter of OneJumpZeroJump, and $\mu$ denotes the population size. Subsequently, Zheng and Doerr Zheng and Doerr (2024b) extended SMS-EMOA to solve a many-objective problem, mmOneJumpZeroJump, showing that the same SPU can bring an acceleration of $\Theta(2^{k}/\mu)=\Theta(2^{k}/(2n/m-2k+3)^{m/2})$ as well, where $m$ is the number of objectives. These works echoed the empirical studies that show the benefit of considering non-elitism in MOEAs Tanabe and Ishibuchi (2019); Liang et al. (2023a).
SPU in Bian et al. (2025); Zheng and Doerr (2024b) introduces randomness by randomly selecting a proportion $p_{s}$ of the combined set of the current population and offspring solution(s) to be directly preserved into the next generation. This essentially gives a chance for inferior solutions to survive, which enables the evolutionary search to go along inferior regions which may be close to Pareto optimal regions. However, a cost of this method is that there is less space for the best solutions in the population. Some very best solutions to the problem (e.g., globally non-dominated solutions), even found by an MOEA, may be discarded during the population update process. This necessitates a large population used. Unfortunately, when the population size is large, the benefit of using SPU may vanish. This is because the benefit of SPU comes from the operation on inferior solutions while the large population size will lead to a small probability of selecting these solutions. For example, for SMS-EMOA solving OneJumpZeroJump, the acceleration of $\Omega(2^{k/2}/(\sqrt{k}\mu^{2}))$ brought by SPU Bian et al. (2025) will vanish when the population size $\mu$ is exponential w.r.t. $k$, e.g., $k=\log n$; for SMS-EMOA solving mmOneJumpZeroJump, the acceleration of $\Theta(2^{k}/(2n/m-2k+3)^{m/2})$ Zheng and Doerr (2024b) will vanish when the number $m$ of objectives is large, e.g., $m\geq k$, because the population size $\mu=(2n/m-2k+3)^{m/2}$ increases rapidly with $m$. This dilemma impacts both the effectiveness and practicality of using SPU in MOEAs.
@@
  • We theoretically show that incorporating an archive mechanism with SPU can reduce the upper bound on the expected running time (even exponentially). For example, comparing Theorems 1 and 3 in Table 1, the expected running time of SMS-EMOA with SPU for solving OneJumpZeroJump, no matter whether an archive is used or not, is $O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$, where $C=e\mu/(p_{s}(1-p_{c}))$. The key difference is that using an archive allows a constant population size, resulting in a significantly smaller $C$ and thus reducing the upper bound significantly. Note that Bian et al. Bian et al. (2024) recently proved the effectiveness of using an archive (bringing polynomial acceleration) for MOEAs, while our analysis for MOEAs with SPU reveals that even exponential acceleration can be obtained.
@@
  • In addition, our derived running time bounds for MOEAs with SPU in the second column of Table 1 are significantly better than the previously known ones Bian et al. (2025); Zheng and Doerr (2024b). This improvement stems from the analysis method of considering both the number and size of jumps across gap of dominated solutions. The method happens to share similarities with that of proving Lemma 12 in Doerr and Lutzeyer (2024). Moreover, our running time bounds are more general, as we consider variable survival probability $p_{s}$ and crossover probability $p_{c}$.
@@
Finally, we give a brief overview about the running time analysis of MOEAs. Over the last decade, there has been an increasing interest for the evolutionary theory community to study MOEAs. Early theoretical works Laumanns et al. (2004a, b); Neumann (2007); Giel and Lehre (2010); Neumann and Theile (2010); Doerr et al. (2013); Qian et al. (2013, 2016); Bian et al. (2018) mainly focus on analyzing the expected running time of a simple MOEA like GSEMO/SEMO. Recently, researchers have begun to examine practical MOEAs. Huang et al. Huang et al. (2021) investigated MOEA/D, assessing the effectiveness of different decomposition methods. Zheng et al. Zheng et al. (2022) conducted the first theoretical analysis of NSGA-II. Bian et al. Bian et al. (2025) analyzed the running time of SMS-EMOA and showed that SPU can bring acceleration. Moreover, Wietheger and Doerr Wietheger and Doerr (2023) demonstrated that NSGA-III Deb and Jain (2014) exhibits superior performance over NSGA-II in solving the tri-objective problem 33OneMinMax. Ren et al. Ren et al. (2024a) analyzed the running time of SPEA2 on three $m$-objective problems. Some other works on well-established MOEAs include Bian and Qian (2022); Zheng and Doerr (2022, 2024a); Cerf et al. (2023); Dang et al. (2023a, b); Doerr and Qu (2023b, c); Doerr et al. (2024, 2025); Lu et al. (2024); Opris et al. (2024); Opris (2025); Ren et al. (2024b).
@@ 3 Running Time Analysis of SMS-EMOA
In this section, we prove that using an archive can bring (even exponential) speedup for the well-established MOEA, SMS-EMOA Beume et al. (2007), with stochastic population update (SPU) solving the OneJumpZeroJump problem. In Section 3.1, we first introduce SMS-EMOA with SPU, and its expected running time for solving OneJumpZeroJump, which is much tighter than the previous results Bian et al. (2025); Zheng and Doerr (2024b). Then, in Section 3.2, we prove that using an archive can significantly reduce the upper bound on the expected running time. Due to space limitation, all mathematical proofs could only be sketched or had to be omitted completely.
@@ 3.1 SMS-EMOA with SPU
In Bian et al. (2025), the SPU method is introduced and shown to be beneficial for the search of MOEAs. During population updates, SPU randomly selects a proportion $p_{s}$ of the current population and the offspring solution(s) to directly survive into the next generation and the removed part is selected from the rest solutions. This implies that each solution, including the worst solution in the population, has at least a probability $p_{s}$ of surviving to the next generation. Specifically, SPU of SMS-EMOA as presented in Algorithm 3 is used to replace the original Population Update procedure in line 12 of Algorithm 1. Note that $p_{s}$ is set to $1/2$ in Bian et al. (2025), while we consider a general $p_{s}$ here.
@@
The expected running time of SMS-EMOA with SPU for solving OneJumpZeroJump has been proven to be $O(\mu n^{k}\cdot\min\{1,\sqrt{k}\mu/2^{k/2}\})$ Bian et al. (2025), which is better than that, i.e., $O(\mu n^{k})$, of the original SMS-EMOA. Intuitively, by by introducing randomness into the population update, the evolutionary search has a chance to go along inferior regions which are close to Pareto optimal regions, thereby making the search easier. Here, we re-prove a tighter upper bound on the expected running time of SMS-EMOA with SPU for solving OneJumpZeroJump, as shown in Theorem 1. It is also more general, as it considers a survival probability $p_{s}\in[1/(\mu+1),1-o(1/\mu))$, rather than just $p_{s}=1/2$ as in Bian et al. (2025). Note that the running time of EAs is often measured by the number of fitness evaluations.
@@
Since the objective vectors in $F_{I}^{*}$ will always be maintained by Lemma 2, we could use the conclusion of Theorem 1 in Bian et al. (2025), which proves that the expected number of fitness evaluations of the first phase is $O(\mu(n\log n+k^{k}))$, where the term $O(\mu k^{k})$ is the expected number of fitness evaluations for finding one objective vector in $F_{I}^{*}$ when event $E$, i.e., any solution in the initial population has at most (k1)(k-1) or at least (nk+1)(n-k+1) 1-bits, happens. By Chernoff bound and $n-2k=\Omega(n)$, an initial solution has at most (k1)(k-1) or at least (nk+1)(n-k+1) 1-bits with probability $\exp(-\Omega(n))$, implying that event $E$ happens with probability $\exp(-\Omega(n))^{\mu}$. Thus, the term $O(\mu k^{k})$ can actually be omitted, implying that the expected number of fitness evaluations of the first phase is $O(\mu n\log n)$.
@@
Under the same conditions as the previous results $O(\mu n^{k}\cdot\min\{1,\sqrt{k}\mu/2^{k/2}\})$ in Bian et al. (2025) and $O(\mu n^{k}\cdot\min\{1,\mu/2^{k}\})$ in Zheng and Doerr (2024b), where SPU is used with $p_{s}=1/2$ and crossover probability $p_{c}=0$, our bound in Theorem 1 becomes $O(\mu n^{k}\cdot(e\ln(2e\mu)/k)^{k-1})$. Since the bound in Zheng and Doerr (2024b) is tighter than that in Bian et al. (2025), we focus on comparing our result only with Zheng and Doerr (2024b). When $k>2e\ln(2e\mu)$, our bound $O(\mu n^{k}\cdot(e\ln(2e\mu)/k)^{k-1})$ brings an improvement ratio of $\Theta(\mu(k/(2e\ln(2e\mu)))^{k-1})$, which can be exponential when $k$ is large, e.g., $k=n/8$. For a very small range $\log\mu\leq k<2e\ln(2e\mu)$, our bound shows no advantage, which is because to minimize Eq. (6), we choose $M=\lceil(k-1)/\ln C-1\rceil$ to round up to the nearest integer, leading to an over-relaxation. If we instead set $M=1$, the same bound can be obtained. The reason of our better bound is in the second phase of finding the Pareto optimal solution $1^{n}$: 1) We consider $M+1$ “jumps” across the gap between dominated solutions, and account for all possible jump sizes to find $1^{n}$; 2) We select the optimal number of jumps, i.e., $M=\lceil(k-1)/\ln C-1\rceil$. Note that Bian et al. Bian et al. (2025) considered only two fixed-size jumps, and although Zheng and Doerr Zheng and Doerr (2024b) accounted for all possible jump sizes, they also reduced the process to two jumps.
@@ 3.2 An Archive is Provably Helpful
In the last section, we have proved that for SMS-EMOA with SPU solving OneJumpZeroJump, the expected running time is $O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$, where $C=e\mu/(p_{s}(1-p_{c}))$. From the analysis, we find that SPU benefits evolutionary search by exploring inferior regions that are close to Pareto optimal areas, but it also requires a larger population size $\mu\geq(n-2k+4)/(1-p_{s})$ to preserve the Pareto optimal solutions discovered. The greater the randomness introduced by SPU (i.e., the larger the value of $p_{s}$), the larger the population size needed. Note that a large population size may diminish the benefit of SPU, because it will lead to a very small probability of selecting specific dominated solutions for reproduction, which is required by SPU. For example, when the population size $\mu$ is exponential w.r.t. $k$, e.g., $k=\log n$, the improvement by SPU will vanish, compared to the expected running time $O(\mu n^{k})$ without SPU Bian et al. (2025).
@@
The smaller population size allows a larger probability of selecting inferior solutions that are close to Pareto optimal areas for reproduction, thus leading to speedup. When $k$ is limited, using only SPU may not bring acceleration. For example, when $k=e\ln(8en)$, if using SPU with $p_{s}=1/2$, $\mu=2n$, and $p_{c}=1/2$, the expected running time of SMS-EMOA on OneJumpZeroJump is $O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})=O(n^{k+1})$ (where $C=e\mu/(p_{s}(1-p_{c}))=8en$), implying no acceleration compared to $O(\mu n^{k})=O(n^{k+1})$ without SPU Bian et al. (2025). However, by adding an archive and reducing the population size $\mu$ to $5$, while keeping other settings unchanged, the bound reduces to $O(n^{k}\cdot(e\ln(20e)/k)^{k-1})$, implying a superpolynomial reduction in the upper bound as $k=e\ln(8en)$. Note that we set $\mu=5$ to satisfy the condition of Theorem 3, i.e., to make $p_{s}=1/2\in[1/(\mu+1),(\mu-2)/(\mu+1)]$. On the other hand, when $k$ is large, using SPU alone can lead to exponential acceleration, and the addition of an archive can further enhance this acceleration. For example, when $k=n/8$, if using SPU with $p_{s}=1/2$, $\mu=2n$, and $p_{c}=1/2$, the expected running time is $O(\mu n^{k}\cdot(8e\ln(8en)/n)^{n/8-1})$, which implies an exponential reduction in the upper bound compared to $O(\mu n^{k})$ without SPU. By adding an archive and setting $\mu$ to $5$, the bound reduces to $O(n^{k}\cdot(8e\ln(20e)/n)^{n/8-1})$, resulting in an improvement ratio of $\Theta(n(\ln(8en)/\ln(20e))^{n/8-1})$, i.e., an exponential reduction in the upper bound.
We also note that the recent work Bian et al. (2024) has proven that using an archive can provide polynomial acceleration for MOEAs for the first time. For example, for SMS-EMOA solving OneMinMax, the expected running time is $O(\mu n\log n)$ both with and without an archive, but the archive allows for a constant $\mu$, achieving an acceleration of $\Theta(n)$. Our work gives another theoretical evidence for the effectiveness of using an archive, and further shows that superpolynomial or even exponential acceleration can be achieved. In our analysis with SPU, a sequence of $M+1$ consecutive jumps are required, which implies continuously selecting specific solutions for reproduction; using an archive allows a small population size, which increases the selection probability significantly and thus leads to greater acceleration.